OpenStudy (anonymous):

solve the equation: x+2x^1/2-4=0

OpenStudy (anonymous):

let u=x^(1/2) u^2 +2u -4 =0

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i dont know where to go from there

OpenStudy (anonymous):

quadratic formula, or completing the sqaure

OpenStudy (anonymous):

so u^2 + 2u = 4 u^2 +2u +1 = 4+1 (u+1)^2 = 5

OpenStudy (anonymous):

u+1 = +_sqrt(5)

OpenStudy (anonymous):

u = -1 +- sqrt(5)

OpenStudy (anonymous):

so sqrt(x) = -1 + sqrt(5) and sqrt(x) =-1-sqrt(5)

OpenStudy (anonymous):

so the first equation gives x= ( sqrt(5) -1 ) and the second eqn gives no solution ( a sqrt cannot be negative )

OpenStudy (anonymous):

ok i got it now